OpenStudy (darleenie):

The length of a rectangular label is 2.5 cm less than twice the width. The perimeter is 40.6 cm. Find the width.

OpenStudy (imaginaryduck):

17.8 cm

OpenStudy (raffle_snaffle):

what!??!?!?!?!

OpenStudy (imaginaryduck):

idk

OpenStudy (raffle_snaffle):

Explain plz. =)

OpenStudy (imaginaryduck):

ok

OpenStudy (imaginaryduck):

the perimeter is 40.6 so I just doubled the length and subtracted it from the perimeter. After that I divided the difference in half.

OpenStudy (imaginaryduck):

I think i read the problem wrong. Hold on.

OpenStudy (darleenie):

I got 7.6cm

OpenStudy (raffle_snaffle):

P = 2(l + w) P = 2l + 2w 40.6cm = 2(2.5cm - 2w) + 2w

OpenStudy (raffle_snaffle):

Solve for w.

OpenStudy (raffle_snaffle):

2.5 cm less than twice the width = 2.5cm - 2w